This question has been answered incorrectly twice. If you cannot do correctly please do not attempt. Thank you.

In an effort to save money for early retirement, an environmental engineering colleague plans to deposit $1,100 per month, starting one month from now, into a fixed rate account that pays 8% per year compounded quarterly. How much will be in the account at the end of 17 years? At the end of 17 years, the account will be$
Answer #1

Future value of annuity = C*((1+r)t−1) / r)

C = $1,100

r = 8% / 4 = 0.02

t = 17*4 = 68

Future value of annuity = 1,100* (((1+0.02)68−1) / 0.02)

Future value of annuity = 1,100 * 142.21 = $156,433.78